OpenStudy (marcelie):

let f(x) = x^2+48x+2000 find f'(x) determine the slope of the tangent line at x=0

OpenStudy (marcelie):

@Photon336

OpenStudy (photon336):

okay so find the slope at x = 0

OpenStudy (photon336):

\[\frac{ d }{ dx }(x^{2}+48x+2000) = 2x+48 \]

OpenStudy (photon336):

\[|_{x=0}(2x+48) = 48 \]

OpenStudy (photon336):

@marcelie

OpenStudy (marcelie):

yup looks good :p